Janet is 6 years older than her sister Rose. In 10 years time, Rose will just be a third of his father's age. Given that their total age is 81. Find their present age.
Answer by ikleyn(52776)

.
Janet is 6 years older than her sister Rose. In 10 years time, Rose will just be a third of his father's age.
Given that their total age is 81. Find their present age.
~~~~~~~~~~


Let x be the Rose's age in 10 years time from now.

Then the Janet's age in 10 years from now is  (x+6) years, and

     the father's age in 10 years from now will be thrice the Rose's age, i.e. 3x years.


For the total of their ages 10 years from now we have THIS equation

    x + (x+6) + 3x = 81.


Simplify and find x

    5x = 81 - 6

    5x = 75

     x = 75/5 = 15.


So, Rose's age will be 15 years in 10 years from now;

    hence, her present age is 15-10 = 5 years.


Then her sister Janet is  5+6 = 11 years old, and the father is 3*(5+10) = 45 years old.


ANSWER.  Their current age is 5 years (Rose);  11 years (Janet)  and  45 years (the father).


CHECK.  In 10 years the sum of their ages will be  (5+10) + (11+10) + (45+10) = 15+21+55 = 91 years.   ! Correct !

Solved and checked.


---------------

The lesson to learn from my post is THIS:

        Choose the unknown/unknowns by a rational way to facilitate the solution (!)